### v1

# Topological Pathways to Two-Dimensional Quantum Turbulence

Abstract
We present a combined experimental and theoretical investigation of the formation and decay kinetics of vortices in two-dimensional, compressible quantum turbulence. We follow the temporal evolution of a quantum fluid of exciton–polaritons, hybrid light–matter quasiparticles, and measure both phase and modulus of the order parameter in the turbulent regime. Fundamental topological conservation laws require that the formation and annihilation of vortices also involve critical points of the velocity field, namely nodes and saddles. Identifying the simplest mechanisms underlying these processes enables us to develop an effective kinetic model that closely aligns with the experimental observations, and shows that different processes are responsible for vortex number growth and decay. These findings underscore the crucial role played by topological constraints in shaping nonlinear, turbulent evolution of two-dimensional quantum fluids.
Topological and dynamical properties of two-dimensional systems are strongly intertwined. This is true not only in condensed matter setups Qi and Zhang (2011); Culcer et al. (2020); Reichhardt et al. (2022) but also for hydrodynamical systems, be these classical or quantum. In classical fluids the identification of topological critical points proves helpful for classifying flow patterns Lighthill (1966); Perry and Fairlie (1975) and studying two-dimensional spatio-temporal chaos and turbulence Moffatt (2001); Rossi et al. (2006); Ouellette and Gollub (2008); de la Cruz et al. (2014); Smith et al. (2017). As for quantum fluids, the importance of quantization of vorticity has been understood long ago Onsager (1949); Feynman (1955) and vortices indeed play a major role in the route to two-dimensional quantum turbulence Nazarenko and Onorato (2006); White et al. (2014); Neely et al. (2013); Sachkou et al. (2019); Gauthier et al. (2019); Johnstone et al. (2019); Eloy et al. (2021); Baker-Rasooli et al. (2023); Panico et al. (2023a), as they do in the classical context McWilliams (1984); Babiano et al. (1987); Benzi et al. (1987); Brachet et al. (1988); Boffetta and Ecke (2012). In this Letter, we further explore the link between dynamical and topological properties in two-dimensional quantum turbulence. We propose to investigate the temporal properties of the quantum fluid velocity field by a novel strategy. The idea is to devise a minimal model for the quantum flow which complies with global topological constraints, without requiring local knowledge of the spatial dynamics of the system. To achieve this, we derive kinetic equations of formation and annihilation of critical points of the velocity field, and apply the approach to a non-equilibrium exciton-polariton fluid. We show that we can reproduce the experimentally observed rate of creation and annihilation of quantized vortices, thus identifying the elementary mechanisms responsible for the increase in the number of vortices –during the quantum turbulence growth– and for its reduction –during the quantum turbulence decay.
We consider a two-dimensional quantum fluid described by a scalar order parameter of the form $\psi(\vec{r},t)=A(\vec{r},t)\exp\{i\Theta(\vec{r},t)\}$. Here the real functions $A$ ($\geq 0$) and $\Theta$ correspond to the amplitude and phase of the order parameter, respectively, and $\vec{r}=(x,y)$. The velocity field of the fluid is $\vec{v}=(\hbar/m)\vec{\nabla}\Theta$ Feynman (1955). In a two-dimensional setting, two topological indices are associated with any domain $D$ delimited by a close contour $C$, namely, the vorticity $I_{\rm V}$ and the Poincaré index $I_{\rm P}$ Nye et al. (1988)
$I_{\rm V}=\frac{1}{2\pi}\oint_{C}{\rm d}\Theta,\;\;\;\;I_{\rm P}=\frac{1}{2\pi}\oint_{C}{\rm d}\varphi,$ (1)
where $\varphi$ denotes the polar angle of $\vec{v}$. Both indices are zero if there are no singular nor stagnation points inside $D$. They assume nontrivial values when the phase $\Theta$ displays extrema (local maxima or minima), saddles, or essential singularities. The corresponding points are nodes (attractive or repulsive), saddles, and quantum vortices, respectively.
Figure 1: Sketch of the streamline pattern around a (positive) vortex, a saddle, and a node (phase minimum). Darker regions correspond to larger values of the phase $\Theta$ of the order parameter. A vortex is a branch point of the phase, the corresponding branch cut is represented by a discontinuity of the color map in the left plot. Nodes and saddles are stagnation points where $\vec{v}=\vec{0}$.
Figure 1 gives the values of the indices attached to each of these points, which we loosely denote as critical points in the following. The vorticity and Poincaré index attached to a given domain are the sum of the indices of all the critical points it contains.
The co-existence of the three types of critical points presented in Fig. 1 has been explicitly experimentally demonstrated in linear Shvartsman and Freund (1995) and nonlinear Congy et al. (2024) optics. The physical system we examine here involves injecting a high-energy polariton superfluid, and allowing it to expand within a circular potential barrier Panico et al. (2023a). The initial kinetic energy provided to the superfluid induces the creation not only of a dense vortex gas but also of a large number of saddles and nodes. The optical nature of polaritons enables us to measure both the phase and the modulus of the order parameter by interferometric techniques Caputo et al. (2019). As shown in Fig. 2, by analyzing the velocity field, we can track the evolution of hundreds of critical points.
Figure 2: A snapshot of the polariton superfluid phase field, with the measured critical points. The three types of critical points are represented with the same symbols as in Fig. 1. A zoom, with streamlines represented as oriented solid lines, highlights the local flow organization, revealing three saddles, two nodes (one attractive and one repulsive), and a negative vortex.
This method enables us to determine, at each time step, the number of vortices, saddles, and nodes present in the system. We emphasize that the presence of nodes is a unique feature of compressible and non-stationary quantum fluids, such as polariton superfluids. These nodes are indeed observed in our experimental results, and we demonstrate below that they play a crucial role in the onset of turbulence.
Model - We consider the following main mechanisms of creation (or annihilation) of critical points in the flow field: (i) the nodes-to-vortices conversion in which two nodes coalesce and give birth to two vortices and (ii) the saddle-node bifurcation which creates one saddle and one node from scratch. These two processes conserve the vorticity and the Poincaré index; they correspond to well-identified bifurcations whose relevance for a two-dimensional quantum fluid has been validated in Ref. Congy et al. (2024). They can be schematically written as chemical reactions:
[…]
where vortices with positive or negative vorticity are denoted as vortex(+) or vortex(-), respectively. The (positive) quantities $a$, $b$, $c$, and $d$ are the reaction rates, see Eq. (3) below. Mechanism (2a) appeared implicitly in works by Indebetouw Indebetouw (1993) and the Soskin group Soskin et al. (1997), then explicitly in Ref. Soskin et al. (1999). Mechanism (2b) is mentioned by Freund in Ref. Freund (1998). Other mechanisms have been observed Congy et al. (2024) which also conserve both the vorticity and the Poincaré index: a saddle can transform into two saddles plus one node in a pitchfork bifurcation, or also a vortex-antivortex pair and two saddles can appear spontaneously (or coalesce) in a process first identified by Nye, Hajnal, and Hannay Nye et al. (1988) which has been termed the “Bristol mechanism” in Ref. Congy et al. (2024). These reactions have been discarded for simplicity reasons (they involve collisions of a larger number of critical points) and also because much less often observed in a previous experiment and in numerical simulations Congy et al. (2024).
From the modeling (2) we write a kinetic equation inspired by rate equations of elementary chemical reactions:
$\begin{split}&\dfrac{{\rm d}V_{\pm}}{{\rm d}t}=aN^{2}-bV_{+}V_{-},\qquad\dfrac{{\rm d}S}{{\rm d}t}=c-dNS,\\ &\dfrac{{\rm d}N}{{\rm d}t}=-2aN^{2}+2bV_{+}V_{-}+c-dNS,\end{split}$ (3)
where $N(t)$ denotes the number of nodes, $S(t)$ the number of saddles, and $V_{+}(t)$ [$V_{-}(t)$] the number of vortices with positive [negative] vorticity. It results from the values of the topological indices listed in Fig. 1 that the total Poincaré index of the system is $I_{\rm\scriptscriptstyle P}=N+V_{+}+V_{-}-S$. It is easily verified that $I_{\rm\scriptscriptstyle P}$ is preserved by the system (3): this comes as no surprise since the elementary processes (2) both conserve the Poincaré index. Similarly, the conserved total vorticity of the system is $V_{+}-V_{-}$. In the following we make the simplifying assumption that this difference is equal to zero: $V_{+}(t)=V_{-}(t)=V(t)/2$ where $V(t)$ is the total number of vortices. This hypothesis is confirmed by the experimental data (such as displayed in Fig. 3) and is certainly sound in the configuration we consider where typically $V(t)\gg 1$ while no external angular momentum is imparted to the system.
Figure 3: (Top) Comparison of the experimental results for $N(t)$, $V(t)$, and $S(t)$ (circles) with the theoretical predictions (lines). Experimental data are averages of four realisations of the same dynamical regime. For $t\leq t_{\rm c}$, the solid lines have been obtained with the numerical integration of Eqs. (4) with the values $\gamma=0.52$, $N_{0}=170$, and $t_{0}=11$ ps foo . For $t>t_{\rm c}$, the dashed lines correspond to the results of (4), while the solid lines come from the numerical resolution of (6) with $\varepsilon=0.045$. (Bottom) Same as above for the quantities $V(t)/S(t)$ and $N(t)/S(t)$. The value of $t_{\rm c}$ is $56$ ps.
Defining the rescaled quantities $\tau=t/t_{0}$, $n=N/N_{0}$, $v=V/N_{0}$, and $s=S/N_{0}$, with $t_{0}=1/\sqrt{2ac}$ and $N_{0}=\sqrt{c/2a}$, makes it possible to cast the system (3) under the following dimensionless form:
$\begin{split}&\dfrac{{\rm d}v}{{\rm d}\tau}=n^{2}-\alpha v^{2},\qquad\dfrac{{\rm d}s}{{\rm d}\tau}=1-\gamma ns,\\ &\dfrac{{\rm d}n}{{\rm d}\tau}=1-n^{2}-\gamma ns+\alpha v^{2},\end{split}$ (4)
where $\alpha=b/(4a)$ and $\gamma=d/(2a)$  [^1] For completeness, note that in terms of the new parameters, the rates of reaction read $a=1/2N_{0}t_{0}$, $b=2\alpha/N_{0}t_{0}$, $c=N_{0}/t_{0}$, and $d=\gamma/N_{0}t_{0}$..
Results - We consider a turbulent regime of the polariton dynamics in which, after fast expansion of the quantum fluid, the onset of vortex clustering and the emergence of the inverse kinetic energy cascade was evidenced on timescales of a few tens of picoseconds Panico et al. (2023a). The numbers of vortices, saddles, and nodes are displayed as circles in the upper part of Fig. 3. At $t=0$, when the fluid hits the barrier, some critical points are already present, having formed during the fluid’s expansion. The turbulent dynamics is initiated at this moment, which we treat as the initial condition. A low-energy data set, where the onset of turbulence is inhibited by dissipation, is shown in the End Matter along with additional details on the experimental configuration.
Figure 4: Experimental snapshots, taken at times $t=15$ and $16$ ps, showing the formation of a vortex-antivortex pair starting from two nearby nodes, as indicated by Eq. (2a). The streamlines of the velocity field $\vec{v}$ are plotted as oriented solid lines on top of the color-coded phase field representing $\Theta(x,y)$.
Let us first focus on the stage of turbulence growth, during which the numbers of vortices and saddles increase significantly (stage 1 in Fig. 3 ). In this time lag, the nucleation of many new vortices and saddles dominates the temporal evolution. This implies imposing $\alpha=0$: indeed, when $\alpha\neq 0$ the system (4) has a fixed point and the numbers of vortices, saddles, and nodes tend to saturate, which is not what is observed in the experiment. We checked that a nonzero value of $\alpha$ always worsens the agreement of the theoretical curve with data: this confirms that in this stage the incompressible kinetic energy of the system is mostly increasing, as required for the establishment of the inverse cascade of kinetic energy.
It is interesting to discuss the values of the rate of reactions in Eqs. (2). In particular $c/d=N_{0}^{2}/\gamma=6\times 10^{4}\gg 1$, implying that the saddle-node bifurcation is mainly unidirectional: the annihilation of a saddle with a node is much less frequent than their creatio ex nihilo. This indicates that the saddle-node formation mechanism (2b) is the real fuel of the whole process. The nodes-to-vortices reaction (2a) merely transmutes some of the nodes into vortices, but could not be effective on its own. This remark is of significance: the spontaneous creation of uniquely a vortex-antivortex pair being topologically forbidden (it would not conserve the Poincaré index) we are in need of an explanation of the increase of the number $V(t)$ of vortices. In the system we consider, the formation of vortices arises from two saddle-nodes bifurcations (2a) followed by a nodes-to-vortices conversion (2b), ultimately resulting in the formation of two saddles and two vortices. This is the reason why the numbers of saddles and of vortices increase at the same pace, see the upper part of Fig. 3.
The results plotted in the lower panel of Fig. 3 indicate that the total Poincaré index is conserved and small. Indeed in this case $N+V=S$, the two quantities $V/S$ and $N/S$ sum to unity, and a minimum of one should correspond to a maximum of the other, as observed in the lower plot of Fig. 3. This property is model-independent: it is a prerequisite which should be embodied in any kinetic model, but its fulfillment is not a guarantee of accuracy of the model. Experimental results confirm the exact conservation of both $I_{\rm V}$ and $I_{\rm P}$ indices in every realization of the measurements.
It is interesting to observe, during the time evolution of the experiment, instances where two nodes transform into a vortex-antivortex pair according to the mechanism (2a). In Fig. 4, we present two consecutive time frames of the same spatial region, highlighting the evolution of the flow pattern from a configuration containing nodes and saddles to one where two nodes are transformed into a vortex and an antivortex.
The results displayed in Fig. 3 show a striking behavior, namely, a sharp temporal transition between stage 1, characterised by the nonlinear growth of the number of vortices/saddles, and stage 2, characterised by a dramatic decrease of the number of vortices and saddles. However, the number of nodes is not experiencing a similar abrupt modification in the same period of time: this supports a scenario which does not involve nodes, still conserving both $I_{\rm V}$ and $I_{\rm P}$. The so-called Bristol mechanism Nye et al. (1988), described by Eq. (5) below, is a perfect candidate:
[…]
In Fig. 5, the Bristol annihilation mechanism is observed in our experiment during the second stage of the dynamics, when dissipation begins to dominate. Two consecutive frames illustrate the simultaneous annihilation of two saddles and a vortex-antivortex pair within a laminar plane flow, marking the dissipation of incompressible kinetic energy from the system.
In view of the significant decrease of the number of vortices and saddles during stage 2, we consider that the rate of reaction $f$ is zero in Eq. (5). Hence, the process is assumed to be unidirectional  [^2] This unidirectional behavior was already observed in the experiment and the numerical simulations of Congy et al. (2024): In this reference the Bristol mechanism was always inducing the concomitant annihilation of two vortices and two saddles, and never their creatio ex nihilo.. The system (4) accordingly modifies to
$\begin{split}&\dfrac{{\rm d}v}{{\rm d}\tau}=n^{2}-\alpha v^{2}-\varepsilon\,v^{2}s^{2},\quad\dfrac{{\rm d}s}{{\rm d}\tau}=1-\gamma ns-\varepsilon\,v^{2}s^{2},\\ &\dfrac{{\rm d}n}{{\rm d}\tau}=1-n^{2}-\gamma ns+\alpha v^{2},\end{split}$ (6)
where $\varepsilon=\tfrac{1}{2}eN_{0}^{3}t_{0}=ec/(8a^{2})$ is the rescaled rate of annihilation of saddles and vortices. We keep for all the other parameters the values previously determined, and during stage 2 we solve the system (6) with $\varepsilon\neq 0$. The corresponding results are displayed in Fig. 3. The agreement of the theoretical curve with the experimental observation supports the idea that after $t=t_{c}$ the system enters a new regime in which the annihilation mechanism (5) acquires an efficiency it previously did not have.
Figure 5: Experimental snapshots of streamlines and the phase field taken during stage 2 of the experiment ($t=76$ and 77 ps), showing the annihilation of two saddles and a vortex-antivortex pair via the Bristol mechanism (5).
It is interesting to ask the question whether the mechanism of Eq. (5) could have been effective earlier, with a rate of reaction $f\neq 0$ explaining the rapid and concomitant increase of $V$ and $S$ during stage 1. The observation of the behavior of $N$ in the same period makes this hypothesis rather unlikely, since $N$ initially increases and then saturates. This advocates for a saddle-node creation process (2b) which then feeds the nodes-to-vortices one (2a). Only this process can explain (i) the occurrence of extrema of $V/S$ and $N/S$ at short times (lower plot of Fig. 3) and (ii) the saturation of $N$ at a slightly later time (upper plot of the same figure). And indeed, it is not possible to accurately reproduce the experimental data on the basis of mechanisms (2a) and (5) only, or (2b) and (5) only.
Discussion - The onset of a turbulent inverse cascade of kinetic energy implies a temporal growth of the incompressible part of the total kinetic energy in the system. It was suggested in Ref. Panico et al. (2023a) that such energy input is at the expenses of the compressible kinetic energy in the quantum flow. This dynamical observation has its topological counterpart in the time window associated to the turbulence (and vortex clustering) growth, where the numbers of vortices and saddles increase, while the number of nodes stays almost constant (stage 1 in Fig. 3). It is reasonable to think that not all vortices participate in the cascade, since they may not have time to correlate, nevertheless their increase reflects in the growth of the incompressible kinetic energy available for the cascade. When the ratio of compressible to incompressible kinetic energy stops growing, dissipation mechanisms prevail and turbulence starts decaying. This happens in our system with the simultaneous rapid decrease of both the vortices and saddles numbers. Interestingly, at the same time the clustering dynamics stops, as observed in Ref. Panico et al. (2023a). Topological constraints also rule the mechanism of such a decay; a process based on four-vortex interactions  [^3] Note that since in the observed dynamics, saddles and vortices have similar temporal evolutions, the decay due to the Bristol mechanism (5) is effectively equivalent to a four-vortex decay process. previously proposed in Nazarenko and Onorato (2007); Groszek et al. (2016); Karl and Gasenzer (2017); Baggaley and Barenghi (2018); Kanai and Zhang (2024), here finds its origin in topological arguments. In the absence of a turbulent regime, the fate of vortices is different. Indeed in such a case, we physically expect a dynamical equilibrium between vortex creation and annihilation processes, in the presence of random, uncorrelated fluctuations. Our model faithfully describe this process, see End Matter.
Our kinetic model is a minimal one: it is the simplest that complies with topological constraints. It provides a global, averaged description of the system based on phenomenological parameters (the rate coefficients) but is not designed to explain why these parameters assume different values in the turbulent or nonturbulent regimes, nor to predict when turbulence growth halts and why its decay (at $t\geq t_{\rm c}$) is so abrupt. Addressing these phenomena requires to account for vortex clustering, i.e., to deal with spatial correlations within the system.
This focus on spatial correlations is crucial in the study of two-dimensional turbulence: since Polyakov’s pioneering contribution Polyakov (1993) it has been shown that the vorticity domains exhibit the same universal scaling arising in critical percolation theory, in both classical Bernard et al. (2006); Puggioni et al. (2020) and quantum Panico et al. (2023b) fluids in the regime of inverse energy cascade. Broadening the scope of our kinetic approach to set up a microscopic model that integrates these statistical properties would therefore be of great interest. Such a model should account for interactions between critical points (such as vortex clustering) within a framework consistent with the conservation of topological indices.
Acknowledgements.
N. P. acknowledges insightful comments by M. V. Berry and M. R. Dennis. D. B., T. C. and A. L. acknowledge the kind hospitality during the workshop “Turbulence and Vortex dynamics in 2D quantum fluids” at the International Center for Theoretical Sciences (ICTS, Bangalore, India), where part of this work was discussed (code: ICTS/QUFLU2024/2). T. C. and N. P. thank the Isaac Newton Institute for Mathematical Sciences, Cambridge, for support and hospitality during the programme Emergent Phenomena in Nonlinear Dispersive Waves, where the work on this paper was partially undertaken. This work was supported by the Italian Ministry of University and Research (MUR) through the PNRR MUR project: ‘National Quantum Science and Technology Institute’ - NQSTI (PE0000023) and the PNRR MUR project: ‘Integrated Infrastructure Initiative in Photonic and Quantum Sciences’ - I-PHOQS (IR0000016). We acknowledge the support of the Quantum Optical Networks based on Exciton-polaritons - (Q-ONE) funding from the HORIZON-EIC-2022-PATHFINDER CHALLENGES EU programme under grant agreement No. 101115575, and of the Neuromorphic Polariton Accelerator - (PolArt) funding from the Horizon-EIC-2023-Pathfinder Open EU programme under grant agreement No. 101130304. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or European Innovation Council and SMEs Executive Agency (EISMEA). Neither the European Union nor the granting authority can be held responsible for them.
The data that support the findings of this article are available upon reasonable request from the authors.
Appendix A End Matter
A.1 Methods
Figure 6: Measured density (top), phase (middle), and critical points (bottom) of the polariton fluid for the high-energy case. At each time frame, the density is normalized to ensure its maximum value is 1. The dashed red circle represents the position of the confining potential. In the bottom row the white diamonds are nodes, the blue ones are saddles, and the red points are vortices (their signs are omitted for legibility). The time frames, from left to right, correspond to: the initial experimental condition ($t=-20$ ps), when the polariton fluid is still mostly localized in the center of the potential and has yet to fill all the available space; $t=15$ ps after the sudden growth in the number of vortices, when the polariton fluid is flowing back after hitting the boundaries; the switching point at around $t=56$ ps, when vortex growth stops and the Bristol mechanism becomes relevant; and finally, a snapshot near the end of the dynamics at $t=80$ ps.
[…]
The time evolution of the polariton fluid is captured using off-axis digital holography, which involves the interference of the signal with a reference pulse (a sample of the excitation beam) with a variable time delay, enabling the retrieval of both the amplitude and phase of the fluid, as illustrated in Fig. 6. The temporal resolution is of around 1 $\mathrm{ps}$ and the spatial resolution is finer than the estimated healing length of the vortices. Each time frame is obtained by integrating over a large number of pulses, given the pulsed pump’s repetition rate of 80 $\mathrm{MHz}$ and the typical integration time of 1 $\mathrm{ms}$. Despite the averaging, spatial inhomogeneities allows for the observation of coherent vortex dynamics. To ensure statistical significance, the analysis averaged four measurements by translating the sample in-plane to eliminate morphological effects. Different spatial configurations can be observed at different sample locations, yet the statistical properties of the observables remain consistent. For each experimental condition and time frame, we identify vortices and critical points by computing the circulation around each point of the two-dimensional phase map and the corresponding velocity field, respectively, and searching for integer multiples of $2\pi$. From a simple comparison, we can further distinguish which of the critical points are nodes.
A.2 Low-energy data set
The data presented in the main text correspond to a turbulent regime in which a high-energy polariton superfluid is injected against a potential barrier. We consider in this section a data set obtained at relatively low injection energy ($E=0.21$ meV here, instead of 1.20 meV in the main text).
Figure 7: Comparison of the experimental results of the low-energy data set for the numbers $N(t)$ of nodes, $V(t)$ of vortices, and $S(t)$ of saddles (circles) with the numerical integration of Eqs. (4) (solid lines).
The experimental data, presented in Fig. 7, display a tendency to saturate: from $t=60$ ps on, $V(t)$, $S(t)$, and $N(t)$ vary quite slowly. To emulate this behavior, the system (4) should have a fixed point and the value of the parameter $\alpha$ should thus be finite. In this case, if we denote as $n_{\infty},v_{\infty},s_{\infty}$ the coordinates of the fixed point, defining $\beta=1/\sqrt{\alpha}$ we get $v_{\infty}=\beta n_{\infty}$, $s_{\infty}=1/(\gamma n_{\infty})$ and
$n_{\infty}=\frac{I_{{\rm\scriptscriptstyle P}0}+\sqrt{I_{{\rm\scriptscriptstyle P}0}^{2}+4(1+\beta)/\gamma}}{2(1+\beta)}.$ (7)
In this expression $I_{{\rm\scriptscriptstyle P}0}=n+v-s$ is the constant value of the rescaled Poincaré index of the whole system ($I_{{\rm\scriptscriptstyle P}0}=I_{\rm\scriptscriptstyle P}/N_{0}$). In the limit $I_{{\rm\scriptscriptstyle P}0}^{2}\ll 4(1+\beta)/\gamma$  [^4] It can be checked a posteriori that this approximation is legitimate: the total experimental Poincaré index is $I_{\rm\scriptscriptstyle P}=-1\pm 2$ whereas $N_{0}$ is of order $10^{2}$: $I_{{\rm\scriptscriptstyle P}0}^{2}\approx 10^{-4}$. The term $4(1+\beta)/\gamma$ is instead of order unity., formula (7) reads $n_{\infty}=[\gamma(1+\beta)]^{-1/2}$ and implies that $s_{\infty}=(1+\beta)n_{\infty}$. Comparing the values $v_{\infty}/n_{\infty}=\beta$ and $s_{\infty}/n_{\infty}=1+\beta$ with the values $V/N\approx 2$ and $S/N\approx 3$ around $t=60$ ps points to a value $\beta\approx 2$ (i.e., $\alpha=0.25$). We found that the choice $\gamma=1$, $N_{0}=80$, $t_{0}=27$ ps, and $\alpha=0.25$ gives a good account of the data set, see Fig. 7.
The system is not here in a turbulent regime such as the one studied in the main text. The injection energy $E$ is smaller, the number of vortices increases at a lower pace, and the stage of decay (stage 2 in Fig. 3) is not reached within the experimental time window. This interpretation of the different behaviors of the two data sets is corroborated by the following evaluation of orders of magnitude: A simple dimensional argument suggests that the characteristic time $t_{0}$ should scale as $t_{0}\propto E^{-1/2}$. And indeed the characteristic times $t_{0}^{\rm low}$ and $t_{0}^{\rm high}$ for the two sets of data are in a ratio $t_{0}^{\rm low}/t_{0}^{\rm high}=27\,{\rm ps}/11\,{\rm ps}=2.45$ which is consistent with the value $\sqrt{E^{\rm high}/E^{\rm low}}=\sqrt{1.20\,{\rm meV}/0.21\,{\rm meV}}=2.39$.